OpenStudy (anonymous):

The area of a square is 144 square feet. What is the length of one side of the square? A. 12 feet B. 288 feet C. 36 feet D. 72 feet

OpenStudy (whpalmer4):

Area of a square is given by A = s^2 = s*s where s = the length of the side (a square is a special case of a rectangle, for which the area formula is A = l*w)

OpenStudy (whpalmer4):

we know A = 144 144 = s^2 = s*s what number multiplied by itself gives you 144?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

thx

OpenStudy (whpalmer4):

36*36 = 144?